OpenStudy (anonymous):

The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the acceleration due to gravity, and t = time. Solve this equation for g, and use your formula to determine the acceleration due to gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work. Pleaseee help c:

OpenStudy (mrnood):

\[d=\frac{ g t ^{2} }{ 2 }\] First multiply both sides of the equation by 2 The divide both sides of the new equation by t^2

OpenStudy (anonymous):

or u can divide by 0.5t^2 and get g=d/0.5t^2

OpenStudy (anonymous):

So d/0.5t^2=g?

OpenStudy (mrnood):

@AntonioCochrane OR you could let the OP do some work for herself....

OpenStudy (anonymous):

lol.sure

OpenStudy (anonymous):

ookay xD

OpenStudy (anonymous):

ya

OpenStudy (mrnood):

so oyu have the formula for g put the numbers in from the question - you can then calculate g

OpenStudy (anonymous):

um thats it

OpenStudy (anonymous):

g=9.8ft/sec?

OpenStudy (mrnood):

It seems that this is the correct answer to this question 490/0.5*10^2 =490/50 = 9.8. PLEASE NOTE TWO THINGS First the units are ft/s^2 not ft/s SECOND - this is NOT the correct value for g on earth - I believe the questioner has got units wrong. IF the question had said 490m (not ft) then the answer of 9.8 m/s^s IS correct value for g on earth.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

u get it now